# Information Science Principles of Machine Learning: A Causal Chain Meta-Framework Based on Formalized Information Mapping

### v10

@@ 2.1 Conceptual Representations of Machine Learning
Another perspective highlights the central role of model parameters (Alquier, 2024), describing machine learning as the process of learning from an object-label set $(X,Y)$ governed by a probability distribution $P$, to produce a family of predictors $\left\{f_{\theta},\theta\in\Theta\right\}$ parameterized by $\theta\in\Theta$. The loss function $\mathcal{l(}f_{\theta}(X),Y)$ quantifies the performance of each predictor, and the expected risk is defined as $R(\theta)=\mathbb{E}_{(X,Y)\sim P}\mathcal{[l(}f_{\theta}(X),Y)]$. The ultimate goal is to find $\theta^{*}$ such that $R\left(\theta^{*}\right)=\inf_{\theta\in\Theta}R(\theta)$.
@@ 2.2 Model Interpretability
Reference (Glanois et al., 2024) argues that, in the context of reinforcement learning, what is considered understandable and what constitutes a good explanation may differ entirely depending on whether the end-user, system designer, or policymaker is being considered. Based on a definition of comprehensibility encompassing simulatability, decomposability, and algorithmic transparency, the authors emphasize that interpretability in reinforcement learning relies on three pillars: interpretable inputs, interpretable models, and interpretable decisions. They further analyze the specific mechanisms for achieving interpretability in each of these aspects.
@@ 2.3 Ethical Safety Assurance
Concerns regarding the ethical safety of AI have a long history. As noted in [10], the proliferation of intelligent and autonomous systems has given rise to significant legal and moral questions, the answers to which impact both producers and consumers of AI technologies. These issues span law, public policy, professional ethics, and philosophical morality, necessitating the expertise of computer scientists, legal scholars, political scientists, and ethicists. While mathematical tools such as formal logic, probability theory, and decision theory have provided important foundations for reasoning and decision-making, numerous challenges remain unresolved.
@@
Building on traditional concepts of assurance, [11] offers a definition of ethical assurance for AI: it is a process that employs structured argumentation to provide auditable guarantees—specifically, that a set of normative claims regarding a system’s ethical properties are justified in light of the available evidence. Accordingly, the general structure of moral assurance is characterized as an iterative process of reflection, action, and adjustment, repeated throughout the design, development, and deployment lifecycle.
@@
In the context of AI/ML-supported healthcare, ethical safety assurance faces more concrete challenges. Reference [12] provides a comprehensive exploration of the multifaceted ethical considerations in deploying AI in healthcare. The authors advocate for multiple strategies, including the implementation of rigorous data security protocols, prioritization of data minimization, increasing diversity in training datasets, use of data augmentation to address imbalance, application of algorithmic fairness techniques to detect and mitigate model bias, regular auditing of learning algorithms, enhancing the interpretability of algorithmic outputs, regulatory oversight based on standardized performance metrics and stringent testing protocols, and the augmentation—rather than replacement—of clinicians’ capabilities as means to achieve ethical safety.
@@
[13] further investigates safety assurance guidelines for machine learning in specific domains such as autonomous systems, establishing a causal chain from ML safety to software safety and ultimately to system safety. The authors propose that ML safety requirements should be derived from the system-level safety requirements, and that these should be evaluated in light of the allocated safety needs, system and software architecture, and operational environment. This framework then serves as the basis for developing safety assurance specifications across all stages of the ML lifecycle, including data preparation, model training, model validation, and model deployment.
@@
Reference [14] conducts an extensive study of the technology practitioners who have the greatest influence on AI ethical safety, proposing a classification framework that encompasses practitioner awareness, perceptions, needs, challenges, and approaches. The study highlights that, in the absence of a universally adopted set of ethical principles for AI, incorporating morality into AI is fundamentally challenged by the lack of clear moral knowledge and the ambiguity of ethical principles.
@@ 2.4 Model Error Estimation
Model performance is fundamentally reflected in the errors it may incur on specific tasks. Reference (Vapnik, 1999) provides key theorems on the consistency of empirical risk minimization (ERM); it characterizes the conditions for both two-sided and one-sided uniform convergence from empirical risk to true risk using VC entropy, VC dimension, and growth functions. Based on VC dimension, it further establishes distribution-independent risk upper bounds and sufficient conditions for fast convergence under both bounded and unbounded loss in finite-sample settings. By controlling the trade-off between VC dimension capacity and empirical risk through nested hypothesis families, it proves the strong consistency and convergence rates of structural risk minimization (SRM). These theoretical foundations have led to practical implementations such as support vector machines (SVM), margin maximization, and kernel methods, offering explanations for their generalization mechanisms.
Reference (Rudin, 2019) replaces global complexity with local empirical Rademacher averages, measuring complexity on subsets with small empirical error. It provides finite-sample error upper bounds via the fixed point of sub-root functions, thus achieving optimal or near-optimal convergence rates inversely proportional to the square root of the number of training samples. Under controlled variance conditions, it presents a unified framework for error estimation, including both distribution-dependent bounds determined by local Rademacher complexity and data-dependent bounds constructed from empirical local Rademacher averages. These results are extended to typical classification and regression tasks, providing computable upper bounds and iterative algorithms for quickly approximating the fixed point.
@@
Reference (Shwartz-Ziv and Tishby, 2017) provides a foundational theoretical framework for meta-reinforcement learning, focusing on in-distribution cross-task generalization and training convergence. It introduces a task distribution T and task-level MDPs, defines generalization error as the performance gap between training and unseen tasks under a meta-policy, and establishes generalization upper bounds determined by the number of training tasks and task complexity, using empirical process theory and concentration inequalities. The analysis further discusses the impact of algorithmic capacity and learning rates on error bounds. Within the non-convex SGD framework, sufficient conditions and rates for finite-step convergence of meta-learning loss are presented, along with guidelines for long-term behavior and hyperparameter selection based on linear and sub-linear types. The theoretical results are validated and the observability of convergence rates is demonstrated through controllable task generation and baseline meta-algorithms.
Reference (Xu et al., 2022) addresses the separation and convergence complexity of continuous and discrete errors in deterministic sampling of generative models. From both theoretical and numerical perspectives, it investigates the convergence properties of deterministic samplers based on probability flow ordinary differential equations (ODEs). Assuming access to score function estimates with $L_{2}$ accuracy, it provides upper bounds on TVD between target and generated distributions in continuous time. For the practical implementation using Runge-Kutta integrators, discrete-time error upper bounds are established. Numerical experiments on 128-dimensional problems are conducted to validate these theoretical error bounds.
@@ 3.1 Definition of Information
Reference (Xu et al., 2014), based on four fundamental postulates, proposes that information is an enabling mapping from the state of an ontology to the state of a carrier, but does not further elaborate on the notions of “state” or “enabling mapping.” Appendix A establishes a higher-order formal system $\mathcal{L}$ and, drawing on general common sense, formulates basic postulates regarding states. It is then proven that any set of states can be represented as the interpretation of a set of WFFs in $\mathcal{L}$ (Xu, 2024), indicating that both ontological states and carrier states correspond to interpretations of sets of WFFs over specific domains of objects and times. Here, we proceed to further clarify the concept of enabling mapping.
@@
Definition 2 (Mathematical Definition of Information) Let $\mathbb{O}$ denote the set of objective entities, $\mathbb{S}$ the set of subjective entities, and $\mathbb{T}$ the set of times, with elements of $\mathbb{O}$, $\mathbb{S}$, and $\mathbb{T}$ partitioned as needed depending on the domain. The ontology $o\mathbb{\subseteq O}\cup\mathbb{S}$, time of occurrence $T_{h}\mathbb{\subseteq T}$, ontological state set $S_{o}\left(o,T_{h}\right)$, carrier $c\mathbb{\subseteq O}$, reflection time $T_{m}\mathbb{\subseteq T}$, and carrier reflection set $S_{c}\left(c,T_{m}\right)$ are all non-empty sets. Then, information $I$ is the enabling mapping from $S_{o}\left(o,T_{h}\right)$ to $S_{c}\left(c,T_{m}\right)$, succinctly denoted as $I=\left\langle o,T_{h},S_{c},c,T_{m},S_{o}\right\rangle$. The set of all information $I$ is called the information space, denoted by $\mathfrak{T}$, which, together with matter and energy, forms the three fundamental constituents of the objective world (Xu et al., 2025).
@@
Reference (Shannon, 1948) is regarded as a foundational work in information theory. The paper proposed that ” $H=-\sum{p_{i}\log p_{i}}$ plays a central role as the measure of information in information theory.” The discussion in this paper includes both the source and channel of information. The source $Is$ is the origin of the information, and it generates events $e_{1}$, $e_{2},$…, $e_{n}$ with probabilities $p_{1}$, $p_{2},$…, $p_{n}$, respectively. The channel $Ic$ is the medium through which information is transmitted. After an event $e_{i}$ occurs at time $T_{h}$ in the source $Is$, the channel $Ic$ receives and transmits the corresponding code $c_{i}$ in the time interval $(T_{h},T_{h}+\tau)$ (where $1\leq i\leq n,\ \tau>0$). We define the ternary predicate ${IsEvent}^{3}(Is,T_{h},e_{i})$ to represent that event $e_{i}$ happened in source $Is$ at time $T_{h}$, and the ternary predicate ${IcCode}^{3}(Ic,T_{m},c_{i})$ to represent that the channel $Ic$ received and transmitted code $c_{i}$ in the time interval $T_{m}$. Then,
@@ 3.3 Information Chain
Reference (Xu et al., 2014) discusses the transmissibility of information through composite mappings. Based on Definitions 1 and 2, the fundamental laws governing information transmission can be characterized more rigorously as follows:
@@ 4.1 Construction of Initial Information
$\begin{split}{\varphi o}_{3}=&{State}^{3}\left(M_{D},t_{1},\Phi^{2}(M_{D},{To}_{h})\right)\land{Input}^{3}(M_{D},t_{1},S_{M_{I}}\left(M_{I},t_{1}\right))\\ &\land{Processable}^{2}(M,S_{M_{I}}\left(M_{I},t_{1}\right))\rightarrow\exists f_{process}^{2}(M,S_{M_{I}}\left(M_{I},t_{1}\right))\\ &({{Eq}^{2}\left(S_{M_{O}}\left(M_{O},t_{2}\right),f_{process}^{2}(M,S_{M_{I}}\left(M_{I},t_{1}\right))\right)\land Output}^{3}(M,t_{2},S_{M_{O}}\left(M_{O},t_{2}\right)))\end{split}$ (4.4)
@@
It should be noted that, since $M_{D}$, $\Phi^{2}(M_{D},{To}_{h})$, and ${State}^{3}\left(M_{D},t_{1},\Phi^{2}(M_{D},{To}_{h})\right)$ all represent sets, ${\varphi o}_{1},{\varphi o}_{2},{\varphi o}_{3}$ may be considered as three well-formed formulas, or—when decomposed with respect to the elements of each set—as three sets of more granular well-formed formulas. Regardless of the situation, both cases satisfy Definition 4 concerning finite automaton states. Therefore, subsequent discussions will encompass both cases without further elaboration.
@@ 4.3 Formal Modeling and Implementation of Model Learning
$\begin{split}{\varphi u}_{3}=&{State}^{3}\left(M,t_{2},\Phi^{2}(M,t_{2})\right)\land{Input}^{3}(M,t_{2},S_{M_{I}}\left(M_{I},t_{2}\right))\\ &\land{Processable}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right))\rightarrow\exists f_{process}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right))\\ &({Eq}^{2}(S_{M_{O}}\left(M_{O},t_{3}\right),f_{process}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right)))\land{Output}^{3}(M,t_{3},S_{M_{O}}\left(M_{O},t_{3}\right))))\end{split}$ (4.14)
@@ 4.6 Multi-Stage Information Causal Chains
Figure 2: Causal Chains of Multiple Information Mappings in Machine Learning.
In machine learning, the four distinct stages—model construction, learning and training, problem input, and answer output—each encompass four levels: conceptual ideas, mind maps, program code, and physical models. Generally, the content within the same level shares a common semantic logic and can thus be represented using a unified formal system. Furthermore, owing to the transmissibility of information, the innermost physical model can be regarded as the result of an enabling mapping from the outer conceptual ideas or mind maps. The output stage can, in turn, be viewed as a reverse causal transmission from the physical model back to conceptual ideas. This demonstrates that the system comprises multiple stages, with each stage containing various combinatorial information causal chains.
@@ 5.1 Theorem of Model Interpretability
Let
$S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))$ $=E_{M},$
$S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i}))$ $=E_{M_{I}},$
$S_{or}(or(t_{i+1}),{Tr}_{h}(t_{i+1}))$ $=E_{M_{O}}$
Then, according to Definition 6, there exist information objects
$Iu(t_{i})$ $=\langle ou(t_{i}),{Tu}_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\rangle,$
$Iq(t_{i})$ $=\langle oq(t_{i}),{Tq}_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\rangle,$
$Ir(t_{i+1})$ $=\langle or(t_{i+1}),{Tr}_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\rangle$
such that
$S_{or}(or(t_{i+1}),{Tr}_{h}(t_{i+1}))$
$\quad\subseteq Cn(S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i})))$
Thus, the necessity of the first part of the theorem is established.
On the other hand, for any $1\leq i\leq n-1$, suppose there exist information objects $Iu(t_{i})=\left\langle ou(t_{i}),{Tu}_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\right\rangle$, $Iq(t_{i})=\left\langle oq(t_{i}),{Tq}_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\right\rangle$, and $Ir(t_{i+1})=\left\langle or(t_{i+1}),{Tr}_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\right\rangle$, such that in the formal system $\mathcal{L}$,
$S_{or}(or(t_{i+1}),{Tr}_{h}(t_{i+1}))$
$\quad\subseteq Cn(S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i})))$
Then, according to Lemma A.1, $E_{M}=S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))$, $E_{M_{I}}=S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i}))$, and $E_{M_{O}}=S_{or}\left(or\left(t_{i+1}\right),{Tr}_{h}\left(t_{i+1}\right)\right)$ are interpretations of $\mathcal{L}$ over the three finite domains $ou(t_{i})\times{Tu}_{h}(t_{i})$, $oq(t_{i})\times{Tq}_{h}(t_{i})$, and $r(t_{i+1})\times{Tr}_{h}(t_{i+1})$, respectively, and $E_{M_{O}}\subseteq Cn(E_{M}\cup E_{M_{I}})$. At the same time, $S_{M}\left(M,t_{i}\right)$, $S_{M_{I}}\left(M_{I},t_{i}\right)$, and $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ are the enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, respectively. By Definition 4, the output $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ of model $M$ under the input $S_{M_{I}}\left(M_{I},t_{i}\right)$ is interpretable in the formal system $\mathcal{L}$. Therefore, the sufficiency of the first part of the theorem is established.

### v11

@@ 2.1 Conceptual Representations of Machine Learning
Another perspective highlights the central role of model parameters (Alquier, 2024), describing machine learning as the process of learning from an object-label set $(X,Y)$ governed by a probability distribution $P$, to produce a family of predictors $\left\{f_{\theta},\theta\in\Theta\right\}$ parameterized by $\theta\in\Theta$. The loss function $L(f_{\theta}(X),Y)$ quantifies the performance of each predictor, and the expected risk is defined as $R(\theta)=\mathbb{E}_{(X,Y)\sim P}\ [L(f_{\theta}(X),Y)]$. The ultimate goal is to find $\theta^{*}$ such that $R\left(\theta^{*}\right)=\inf_{\theta\in\Theta}R(\theta)$.
@@ 2.2 Model Interpretability
(Glanois et al., 2024) argues that, in the context of reinforcement learning, what is considered understandable and what constitutes a good explanation may differ entirely depending on whether the end-user, system designer, or policymaker is being considered. Based on a definition of comprehensibility encompassing simulatability, decomposability, and algorithmic transparency, the authors emphasize that interpretability in reinforcement learning relies on three pillars: interpretable inputs, interpretable models, and interpretable decisions. They further analyze the specific mechanisms for achieving interpretability in each of these aspects.
@@ 2.3 Ethical Safety Assurance
Concerns regarding the ethical safety of AI have a long history. As noted in (Russell et al., 2015), the proliferation of intelligent and autonomous systems has given rise to significant legal and moral questions, the answers to which impact both producers and consumers of AI technologies. These issues span law, public policy, professional ethics, and philosophical morality, necessitating the expertise of computer scientists, legal scholars, political scientists, and ethicists. While mathematical tools such as formal logic, probability theory, and decision theory have provided important foundations for reasoning and decision-making, numerous challenges remain unresolved.
@@
Building on traditional concepts of assurance, (Burr and Leslie, 2023) offers a definition of ethical assurance for AI: it is a process that employs structured argumentation to provide auditable guarantees—specifically, that a set of normative claims regarding a system’s ethical properties are justified in light of the available evidence. Accordingly, the general structure of moral assurance is characterized as an iterative process of reflection, action, and adjustment, repeated throughout the design, development, and deployment lifecycle.
@@
In the context of AI/ML-supported healthcare, ethical safety assurance faces more concrete challenges. (Tilala et al., 2024) provides a comprehensive exploration of the multifaceted ethical considerations in deploying AI in healthcare. The authors advocate for multiple strategies, including the implementation of rigorous data security protocols, prioritization of data minimization, increasing diversity in training datasets, use of data augmentation to address imbalance, application of algorithmic fairness techniques to detect and mitigate model bias, regular auditing of learning algorithms, enhancing the interpretability of algorithmic outputs, regulatory oversight based on standardized performance metrics and stringent testing protocols, and the augmentation—rather than replacement—of clinicians’ capabilities as means to achieve ethical safety.
@@
(Hawkins et al., 2021) further investigates safety assurance guidelines for machine learning in specific domains such as autonomous systems, establishing a causal chain from ML safety to software safety and ultimately to system safety. The authors propose that ML safety requirements should be derived from the system-level safety requirements, and that these should be evaluated in light of the allocated safety needs, system and software architecture, and operational environment. This framework then serves as the basis for developing safety assurance specifications across all stages of the ML lifecycle, including data preparation, model training, model validation, and model deployment.
@@
(Pant et al., 2024) conducts an extensive study of the technology practitioners who have the greatest influence on AI ethical safety, proposing a classification framework that encompasses practitioner awareness, perceptions, needs, challenges, and approaches. The study highlights that, in the absence of a universally adopted set of ethical principles for AI, incorporating morality into AI is fundamentally challenged by the lack of clear moral knowledge and the ambiguity of ethical principles.
@@ 2.4 Model Error Estimation
Model performance is fundamentally reflected in the errors it may incur on specific tasks. (Vapnik, 1999) provides key theorems on the consistency of empirical risk minimization (ERM); it characterizes the conditions for both two-sided and one-sided uniform convergence from empirical risk to true risk using VC entropy, VC dimension, and growth functions. Based on VC dimension, it further establishes distribution-independent risk upper bounds and sufficient conditions for fast convergence under both bounded and unbounded loss in finite-sample settings. By controlling the trade-off between VC dimension capacity and empirical risk through nested hypothesis families, it proves the strong consistency and convergence rates of structural risk minimization (SRM). These theoretical foundations have led to practical implementations such as support vector machines (SVM), margin maximization, and kernel methods, offering explanations for their generalization mechanisms.
(Rudin, 2019) replaces global complexity with local empirical Rademacher averages, measuring complexity on subsets with small empirical error. It provides finite-sample error upper bounds via the fixed point of sub-root functions, thus achieving optimal or near-optimal convergence rates inversely proportional to the square root of the number of training samples. Under controlled variance conditions, it presents a unified framework for error estimation, including both distribution-dependent bounds determined by local Rademacher complexity and data-dependent bounds constructed from empirical local Rademacher averages. These results are extended to typical classification and regression tasks, providing computable upper bounds and iterative algorithms for quickly approximating the fixed point.
@@
(Shwartz-Ziv and Tishby, 2017) provides a foundational theoretical framework for meta-reinforcement learning, focusing on in-distribution cross-task generalization and training convergence. It introduces a task distribution T and task-level MDPs, defines generalization error as the performance gap between training and unseen tasks under a meta-policy, and establishes generalization upper bounds determined by the number of training tasks and task complexity, using empirical process theory and concentration inequalities. The analysis further discusses the impact of algorithmic capacity and learning rates on error bounds. Within the non-convex SGD framework, sufficient conditions and rates for finite-step convergence of meta-learning loss are presented, along with guidelines for long-term behavior and hyperparameter selection based on linear and sub-linear types. The theoretical results are validated and the observability of convergence rates is demonstrated through controllable task generation and baseline meta-algorithms.
(Xu et al., 2022) addresses the separation and convergence complexity of continuous and discrete errors in deterministic sampling of generative models. From both theoretical and numerical perspectives, it investigates the convergence properties of deterministic samplers based on probability flow ordinary differential equations (ODEs). Assuming access to score function estimates with $L_{2}$ accuracy, it provides upper bounds on TVD between target and generated distributions in continuous time. For the practical implementation using Runge-Kutta integrators, discrete-time error upper bounds are established. Numerical experiments on 128-dimensional problems are conducted to validate these theoretical error bounds.
@@ 3.1 Definition of Information
(Xu et al., 2014), based on four fundamental postulates, proposes that information is an enabling mapping from the state of an ontology to the state of a carrier, but does not further elaborate on the notions of “state” or “enabling mapping.” Appendix A establishes a higher-order formal system $\mathcal{L}$ and, drawing on general common sense, formulates basic postulates regarding states. It is then proven that any set of states can be represented as the interpretation of a set of WFFs in $\mathcal{L}$ (Qiu and Xu, 2025), indicating that both ontological states and carrier states correspond to interpretations of sets of WFFs over specific domains of objects and times. Here, we proceed to further clarify the concept of enabling mapping.
@@
Definition 2 (Mathematical Definition of Information) Let $\mathbb{O}$ denote the set of objective entities, $\mathbb{S}$ the set of subjective entities, and $\mathbb{T}$ the set of times, with elements of $\mathbb{O}$, $\mathbb{S}$, and $\mathbb{T}$ partitioned as needed depending on the domain. The ontology $o\mathbb{\subseteq O}\cup\mathbb{S}$, time of occurrence $T_{h}\mathbb{\subseteq T}$, ontological state set $S_{o}\left(o,T_{h}\right)$, carrier $c\mathbb{\subseteq O}$, reflection time $T_{m}\mathbb{\subseteq T}$, and carrier reflection set $S_{c}\left(c,T_{m}\right)$ are all non-empty sets. Then, information $I$ is the enabling mapping from $S_{o}\left(o,T_{h}\right)$ to $S_{c}\left(c,T_{m}\right)$, succinctly denoted as $I=\left\langle o,T_{h},S_{c},c,T_{m},S_{o}\right\rangle$. The set of all information $I$ is called the information space, denoted by $\mathfrak{T}$, which, together with matter and energy, forms the three fundamental constituents of the objective world (Wiener, 2019).
@@
(Shannon, 1948) is regarded as a foundational work in information theory. The paper proposed that ” $H=-\sum{p_{i}\log p_{i}}$ plays a central role as the measure of information in information theory.” The discussion in this paper includes both the source and channel of information. The source $Is$ is the origin of the information, and it generates events $e_{1}$, $e_{2},$…, $e_{n}$ with probabilities $p_{1}$, $p_{2},$…, $p_{n}$, respectively. The channel $Ic$ is the medium through which information is transmitted. After an event $e_{i}$ occurs at time $T_{h}$ in the source $Is$, the channel $Ic$ receives and transmits the corresponding code $c_{i}$ in the time interval $(T_{h},T_{h}+\tau)$ (where $1\leq i\leq n,\ \tau>0$). We define the ternary predicate ${IsEvent}^{3}(Is,T_{h},e_{i})$ to represent that event $e_{i}$ happened in source $Is$ at time $T_{h}$, and the ternary predicate ${IcCode}^{3}(Ic,T_{m},c_{i})$ to represent that the channel $Ic$ received and transmitted code $c_{i}$ in the time interval $T_{m}$. Then,
@@ 3.3 Information Chain
(Xu et al., 2014) discusses the transmissibility of information through composite mappings. Based on Definitions 1 and 2, the fundamental laws governing information transmission can be characterized more rigorously as follows:
@@ 4.1 Construction of Initial Information
$\varphi_{O_{3}}=$ $State^{3}(M_{D},t_{1},\Phi^{2}(M_{D},To_{h}))\wedge Input^{3}(M_{D},t_{1},S_{M},(M_{I},t_{1}))$ (4.4)
$\wedge Processable^{2}(M,S_{M},(M_{I},t_{1}))\rightarrow\exists f^{2}_{process}(M,S_{M},(M_{I},t_{1}))$
$(Eq^{2}(S_{M_{0}}(M_{0},t_{2}),f^{2}_{process}(M,S_{M},(M_{I},t_{1}))))\wedge Output^{3}(M,t_{2},S_{M_{0}}(M_{0},t_{2}))$
@@
It should be noted that, since $M_{D}$, $\Phi^{2}(M_{D},{To}_{h})$, and ${State}^{3}\left(M_{D},t_{1},\Phi^{2}(M_{D},{To}_{h})\right)$ all represent sets, ${\varphi o}_{1},{\varphi o}_{2},{\varphi o}_{3}$ may be considered as three well-formed formulas, or—when decomposed with respect to the elements of each set—as three sets of more granular well-formed formulas. Regardless of the situation, both cases satisfy Definition B.1 concerning finite automaton states. Therefore, subsequent discussions will encompass both cases without further elaboration.
@@ 4.3 Formal Modeling and Implementation of Model Learning
${\varphi u}_{3}=$ $State^{3}\left(M,t_{2},\Phi^{2}(M,t_{2})\right)\wedge Input^{3}\left(M,t_{2},S_{M_{I}}(M_{I},t_{2})\right)$ (4.14)
$\wedge Processable^{2}\left(M,S_{M_{I}}(M_{I},t_{2})\right)\rightarrow\exists f_{process}^{2}\left(M,S_{M_{I}}(M_{I},t_{2})\right)$
$\left(Eq^{2}\left(S_{M_{O}}(M_{O},t_{3}),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2}))\right)\wedge Output^{3}\left(M,t_{3},S_{M_{O}}(M_{O},t_{3})\right)\right)$
@@ 4.6 Multi-Stage Information Causal Chains
Figure 2: Causal Chains of Multiple Information Mappings in Machine Learning. In machine learning, the four distinct stages—model construction, learning and training, problem input, and answer output—each encompass four levels: conceptual ideas, mind maps, program code, and physical models. Generally, the content within the same level shares a common semantic logic and can thus be represented using a unified formal system. Furthermore, owing to the transmissibility of information, the innermost physical model can be regarded as the result of an enabling mapping from the outer conceptual ideas or mind maps. The output stage can, in turn, be viewed as a reverse causal transmission from the physical model back to conceptual ideas. This demonstrates that the system comprises multiple stages, with each stage containing various combinatorial information causal chains.
@@ 5.1 Theorem of Model Interpretability
Let $S_{on}(ou(t_{i}),Tu_{h}(t_{i}))=E_{M}$, $S_{op}(oq(t_{i}),Tq_{h}(t_{i}))=E_{M_{I}}$ and $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))=E_{M_{O}}$. Then, according to Definition 4, there exist information objects $Ig(t_{i})=\langle oq(t_{i}),Tq_{h}(t_{i}),S_{op},M_{I},t_{i},S_{M_{I}}\rangle$ and $Ir(t_{i+1})=\langle or(t_{i+1}),Tr_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\rangle$ such that $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\subseteq Cn(S_{on}(ou(t_{i}),Tu_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),Tq_{h}(t_{i})))$. Thus, the necessity of the first part of the theorem is established.
On the other hand, for any $1\leq i\leq n-1$, suppose there exist information objects $Iu(t_{i})=\left\langle ou(t_{i}),{Tu}_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\right\rangle$, $Iq(t_{i})=\left\langle oq(t_{i}),{Tq}_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\right\rangle$, and $Ir(t_{i+1})=\left\langle or(t_{i+1}),{Tr}_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\right\rangle$, such that in the formal system $\mathcal{L}$, $S_{or}(or(t_{i+1}),{Tr}_{h}(t_{i+1}))\quad\subseteq Cn(S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i})))$. Then, according to Lemma A.1, $E_{M}=S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))$, $E_{M_{I}}=S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i}))$, and $E_{M_{O}}=S_{or}\left(or\left(t_{i+1}\right),{Tr}_{h}\left(t_{i+1}\right)\right)$ are interpretations of $\mathcal{L}$ over the three finite domains $ou(t_{i})\times{Tu}_{h}(t_{i})$, $oq(t_{i})\times{Tq}_{h}(t_{i})$, and $r(t_{i+1})\times{Tr}_{h}(t_{i+1})$, respectively, and $E_{M_{O}}\subseteq Cn(E_{M}\cup E_{M_{I}})$. At the same time, $S_{M}\left(M,t_{i}\right)$, $S_{M_{I}}\left(M_{I},t_{i}\right)$, and $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ are the enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, respectively. By Definition 4, the output $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ of model $M$ under the input $S_{M_{I}}\left(M_{I},t_{i}\right)$ is interpretable in the formal system $\mathcal{L}$. Therefore, the sufficiency of the first part of the theorem is established.